OpenStudy (anonymous):

Solve the follow equation in Radians to two decimal places. 3cos x/2 +1 = 0

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

Is this your problem?\[3\cos(\frac{ x }{ 2 })+1=0\]

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

@BlixFace

OpenStudy (anonymous):

The first step is to move everything that you can to the right side. \[\cos \left( \frac{ x }{ 2 } \right)=-\frac{ 1 }{ 3 }\]Now we find where inverse cosine is -1/3, which is 1.9106 \[\ \left( \frac{ x }{ 2 } \right)=1.9106+2k \pi\] Next step is to get rid of the 2 inside the parenthesis. \[x=2(1.9106+2k \pi )\rightarrow 3.8212+4k \pi\] \[\theta=3.82+4k \pi\]

OpenStudy (anonymous):

@blixface where does the 2kπ come from?

OpenStudy (anonymous):

It's to express all solutions for cosine. You can try to review your textbook for solving trigonometric equations for a more detailed explanation.